### v3

# Augment then Smooth: Reconciling Differential Privacy with Certified Robustness

Abstract
Machine learning models are susceptible to a variety of attacks that can erode trust, including attacks against the privacy of training data, and adversarial examples that jeopardize model accuracy. Differential privacy and certified robustness are effective frameworks for combating these two threats respectively, as they each provide future-proof guarantees. However, we show that standard differentially private model training is insufficient for providing strong certified robustness guarantees. Indeed, combining differential privacy and certified robustness in a single system is non-trivial, leading previous works to introduce complex training schemes that lack flexibility. In this work, we present DP-CERT, a simple and effective method that achieves both privacy and robustness guarantees simultaneously by integrating randomized smoothing into standard differentially private model training. Compared to the leading prior work, DP-CERT gives up to a 2.5$\times$ increase in certified accuracy for the same differential privacy guarantee on CIFAR10. Through in-depth per-sample metric analysis, we find that larger certifiable radii correlate with smaller local Lipschitz constants, and show that DP-CERT effectively reduces Lipschitz constants compared to other differentially private training methods. Code is available at github.com/layer6ai-labs/dp-cert.
1 Introduction
Machine learning (ML) models are becoming increasingly trusted in critical settings despite an incomplete understanding of their properties. This raises questions about the trustworthiness of those models, encompassing aspects such as privacy, robustness, and more. Society at large might expect all of these aspects to be accounted for simultaneously as ML’s influence on everyday life expands, but scientists and practitioners still mostly grapple with each aspect individually.
We aim to reconcile two key objectives of trustworthy ML, namely privacy and robustness. Privacy in the context of ML manifests as the requirement that a model does not leak information about the data it was trained on (Papernot et al., 2018), such as revealing whether or not certain data points were included in the training dataset (Shokri et al., 2017) or what characteristics they exhibit (Fredrikson et al., 2015). In our study, robustness refers to the requirement that a model’s prediction should not change when its inputs are perturbed at test-time, even in the worst case of adversarially chosen perturbations.
The current gold standard for providing privacy guarantees is differential privacy (DP) (Dwork & Roth, 2014). In ML, DP produces mathematically rigorous privacy guarantees by limiting the impact of each training data point on the final model. This is achieved by clipping per-sample gradients, and adding a well-calibrated amount of noise to all model updates. Clipping limits the sensitivity of the training algorithm, while the addition of noise ensures that training will be more likely to output similar models when any individual data point is added to or removed from the training dataset.
To quantify robustness, we focus on test-time certified robustness (CR) (Wong & Kolter, 2018; Raghunathan et al., 2018), which provides probabilistic guarantees that perturbations of a certain magnitude will not change a model’s prediction, regardless of what attack strategy is used to modify the inputs. A common approach for certifying robustness is randomized smoothing, where a classifier’s outputs are averaged over a distribution surrounding the data point (Lecuyer et al., 2019; Li et al., 2019; Cohen et al., 2019).
Unfortunately, the two aims of providing DP and CR guarantees are in conflict, both empirically and conceptually. The clipping and noise addition used in DP training can impede the convergence of models (Tramèr & Boneh, 2021) and yield decision boundaries that are less smooth (Hayes et al., 2022), negatively impacting robustness (Fawzi et al., 2018). There is also substantial empirical evidence that providing privacy guarantees is in tension with robustness (Boenisch et al., 2021; Tursynbek et al., 2021). Integrating robustness measures into private training remains challenging because most methods to increase robustness use random or adversarial augmentations of training data points, which do not align well with DP training. Conceptually, augmenting an input increases the sensitivity of private training to it, and thereby provides additional avenues for information leakage. From a practical viewpoint, since gradients are computed on a per-example basis for DP, augmentations drastically increase the time and memory costs of training.
Present Work.
We study the possible pitfalls of combining DP and CR. Through our analysis and ablation studies combining randomized smoothing techniques with DP training, we show that standard DP training of ML models is insufficient to provide strong CR results. Surprisingly, we find that prior complex methods used to regularize models to improve CR are unnecessary (Phan et al., 2019; 2020; Tang et al., 2022). Instead, we propose DP-CERT, a straightforward and adaptable framework for integrating CR into standard DP training which effectively incorporates augmentations while managing the additional privacy risks. Compared to the few existing approaches for integrating DP and CR, our proposed DP-CERT method is simpler, more flexible, and does not require extra network components. This decreased complexity is important for the widespread adoption of private and robust ML. DP-CERT even surpasses the state-of-the-art robustness guarantees on CIFAR10 under equivalent privacy guarantees.
To provide some insight into the mechanisms by which regularization methods improve CR for private models, we analyze CR on a per-data point basis. Using the gradient norm, Hessian spectral norm, and local Lipschitz constant, we find that the certifiable radius has a negative log-linear correlation with these quantities, and compare their distributions across methods. We conclude with concrete recommendations of best practices for the community to achieve CR and DP simultaneously.
2 Preliminaries
Problem Setup.
We consider a $Y$-class classification task with a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{1,{...},Y\}$. Let $f_{\theta}:\mathbb{R}^{d}\xrightarrow{}\{1,{...},Y\}$ be a neural network classifier with parameters $\theta$, and $F_{\theta}$ denote the soft classifier which outputs the probability distribution, such that $f_{\theta}(x)=\operatorname*{arg\,max}_{y\in\{1,{...},Y\}}F_{\theta}(x)_{y}$, where $F_{\theta}(x)_{y}$ denotes the model probability of $x$ being a member of class $y$.
Differential Privacy and DPSGD.
We rely on the rigorous framework of differential privacy (DP) (Dwork et al., 2006) to obtain models with privacy guarantees. DP ensures that a model’s weights at the end of training will be similar in distribution whether or not a particular data point was included in the training set. More formally, let $D$ and $D^{\prime}$ be two potential training datasets for a model $f_{\theta}$ that differ in only one data point. The training mechanism $M$ guarantees $(\varepsilon,\delta)$-DP if, for all possible sets of outcomes $S$ of the training process, it holds that $\mathrm{Pr}\left[M(D)\in S\right]\leq e^{\varepsilon}\mathrm{Pr}\left[M(D^{\prime})\in S\right]+\delta$. The parameter $\varepsilon$ specifies the privacy level, with smaller $\varepsilon$ yielding higher privacy, while $\delta$ quantifies the probability of a catastrophic failure of privacy (Kifer et al., 2022).
To obtain a differentially private variant of stochastic gradient descent (SGD), two modifications need to be made (Abadi et al., 2016). First, the individual gradients of each data point are clipped to a norm $C$ to limit the sensitivity of the model update caused by each data point. Second, choosing a noise level $\rho$, noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$ is added to the aggregated gradients to prevent the changes to the model from revealing too much information about individual data points. The resulting algorithm DPSGD is detailed in Algorithm 1. We provide a more thorough introduction to DP and explanation of Algorithm 1 in Appendix A.1.
Algorithm 1 Standard DPSGD, adapted from (Abadi et al., 2016).
0:  Private training set $D=\{(x_{i},y_{i})\mid i\in[N_{\text{prv}}]\}$, loss function $L(\theta_{t};x,y)$, Parameters: learning rate $\lambda_{t}$, noise scale $\rho$, group size $B$, gradient norm bound $C$.
1:  Initialize ${\theta}_{0}$ randomly
2:  for $t\in[T]$ do
3:     Sample mini-batch $B_{t}$ with sampling probability $B/N_{\text{prv}}$Poisson sampling}
4:     For each $i\in B_{t}$, compute ${\mathbf{g}}_{t}(x_{i})\leftarrow\nabla_{\theta}L({\theta}_{t};x_{i},y_{i})$Compute per-sample gradients}
5:     $\bar{{\mathbf{g}}}_{t}(x_{i})\leftarrow{\mathbf{g}}_{t}(x_{i})/\max\left(1,\frac{\|{\mathbf{g}}_{t}(x_{i})\|_{2}}{C}\right)$Clip gradients}
6:     $\tilde{{\mathbf{g}}}_{t}\leftarrow\frac{1}{|B_{t}|}\left(\sum_{i}\bar{{\mathbf{g}}}_{t}(x_{i})+\mathcal{N}\left(0,\rho^{2}C^{2}{\mathbf{I}}\right)\right)$Add noise to aggregated gradient}
7:      ${\theta}_{t+1}\leftarrow{\theta}_{t}-\lambda_{t}\tilde{{\mathbf{g}}}_{t}$Gradient descent}
8:  end for
9:  Output ${\theta}_{T}$ and compute the overall privacy cost $(\varepsilon,\delta)$ using a privacy accounting method.
Certified Robustness.
Adversarial examples are a well-studied phenomenon in ML, in which an input to a model at test-time is perturbed in ways that do not alter its semantics yet cause the model to misclassify the perturbed input (Biggio et al., 2013; Szegedy et al., 2013; Goodfellow et al., 2014). Formally, for a given labeled datapoint $(x,y)$ and classifier $f$, an $(L_{p},\zeta)$-adversary aims to create an adversarial example $x^{\prime}$ such that $\|x^{\prime}-x\|_{p}<\zeta$ and $f(x^{\prime})\neq y$. Despite much research, the most common defense against adversarial examples remains adversarial training (Goodfellow et al., 2014; Zhang et al., 2019). While adversarial training improves robustness to known algorithms for finding adversarial examples, it does not guarantee that a model will be robust to all adversarial examples. This motivates developing certifiable guarantees of robustness which provide a lower bound $r$ on the distance between a correctly classified input and any adversarial example that may be misclassified (Wong & Kolter, 2018; Raghunathan et al., 2018). This lower bound is also known as the certification radius.
Randomized Smoothing.
A popular approach for establishing certified robustness (CR) guarantees is through probabilistic robustness verification which, with high probability, verifies that no adversarial examples exist within a radius $r$ of the original input (Li et al., 2023). The most common method smooths a classifier by averaging the class predictions of $f$ using a smoothing distribution $\mu$ (Lecuyer et al., 2019; Li et al., 2019; Cohen et al., 2019),
[…]
where $\mathbb{I}[a,b]=1{\iff}a=b$ and 00 otherwise (Li et al., 2023). As computing the integral in Equation (1) is intractable, Monte Carlo sampling is used. We denote the approximation of $\hat{g}$ given by Monte Carlo sampling as $g$. Smoothed classifiers are evaluated in terms of their certified accuracy—the fraction of samples correctly classified when certifying robustness at a given radius $r$.
A tight $L_{2}$ radius was obtained by Cohen et al. (2019) when using isotropic Gaussian noise $\mu=\mathcal{N}(x,\sigma^{2}\mathbf{I})$, where $\sigma$ is a hyperparameter that controls a robustness/accuracy tradeoff. In particular, it was proved that for any base classifier $f$, the Gaussian smoothed classifier $g$ is robust around an input $x$ with radius $r=\frac{\sigma}{2}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B}))$ where $p_{A}$ and $p_{B}$ denote the probabilities of $c_{A}$ and $c_{B}$, the most and second-most probable classes returned by $g(x)$, and $\Phi^{-1}$ is the inverse of the standard Gaussian CDF. The exact probabilities $p_{A}$ and $p_{B}$ are not needed; one can use lower $\underline{p_{A}}\leq p_{A}$ and upper $\overline{p_{B}}\geq p_{B}$ bounds instead, approximated by Monte Carlo sampling. The output of the smoothed classifier $g(x)$ is approximated by aggregating the predictions of a base classifier $f(x+\eta)$ for $\eta\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. As a high dimensional standard Gaussian assigns almost no mass near its mean, ensuring that $g(x)$ is accurate at large certification radii $r$ requires $f$ to be accurate on Gaussian perturbed data (Gao et al., 2022).
Related Work.
Previous works combining DP and CR deviate significantly from the standard DPSGD template, and as a result become more difficult to integrate into DP training pipelines.
Phan et al. (2019) introduced Secure-SGD, the first framework aimed at achieving both CR and DP simultaneously. They used a PixelDP approach (Lecuyer et al., 2019) to attain CR and introduced the heterogeneous Gaussian mechanism, which adds heterogeneous noise across elements of the gradient vector.
Phan et al. (2020) introduced StoBatch which employs an autoencoder (Hinton & Zemel, 1993) and a functional DP mechanism, objective perturbation (Zhang et al., 2012), to reconstruct input examples while preserving DP. Subsequently, reconstructed data is used to train a deep neural network, along with adversarial training (Tramèr et al., 2017) to improve robustness.
Tang et al. (2022) proposed perturbing input gradients during training and introduced the multivariate Gaussian mechanism which allows them to achieve the same DP guarantee with less noise added. Their method TransDenoiser follows the architecture of denoised smoothing, adding differentially private noise to a pre-trained classifier.
We note that these prior approaches either add trainable model components increasing the overall complexity, or lack the flexibility to incorporate the latest adversarial training methods.
3 Method
Training ML models to be both differentially private and certifiably robust poses two main challenges. In this section, we describe these challenges and how our DP-CERT framework meets them by effectively improving certified robustness while preserving privacy.
The first challenge surfaces around the use of adversarial training or augmentations with DPSGD. As shown by Cohen et al. (2019), data augmentations used for training can enhance a model’s CR, however, augmented data points could leak private information about the originals. Previous works combining DP and CR have proposed adding noise or adversarial examples during training, but deviate from the standard DPSGD template to address the privacy risks (Phan et al., 2019; 2020; Tang et al., 2022).
The second challenge is that gradient clipping and noise addition in DPSGD harms the convergence rate of training (Chen et al., 2020; Tramèr & Boneh, 2021; Bu et al., 2023a), while restrictive privacy budgets may require stopping training prior to convergence. Robustness on the other hand suffers when models are not converged, as having large gradients at test points makes finding adversarial examples easier (Fawzi et al., 2018). Strategies must be employed to improve the convergence of private training, or regularize gradient size.
Figure 1: The DP-CERT training framework for providing strong CR guarantees within DPSGD.
We aim to make CR feasible within the standard pipeline of DPSGD, with state-of-the-art convergence and proper accounting for additional privacy risks. Our DP-CERT framework effectively trains with augmented samples while preserving privacy, and integrates advancements in adversarial training and regularizers to enhance certifiable robustness (Salman et al., 2019; Li et al., 2019; Zhai et al., 2020). A schematic of DP-CERT is given in Figure 1 and the training framework in described in Algorithm 2 below. We describe the components in turn.
Augmentation Multiplicity.
The first step in Figure 1 involves training on augmented data with the aim of improving robustness. For each data point $(x_{i},y_{i})$, we obtain $K$ augmented data points $(x_{i}^{j},y_{i})$, where $j\in\{1,...,K\}$ and $x_{i}^{j}$ is the $j$-th augmented version of $x_{i}$. For notational convenience, we use $x_{i}^{0}$ to denote the original data point $x_{i}$. To improve CR, augmentations may be generated adversarially, but since this is computationally expensive, we also consider Gaussian perturbations which, as shown by Cohen et al. (2019), also can enhance a model’s certified robustness. When using Gaussian perturbations instead of adversarial training, we define $x_{i}^{j}=x_{i}+\eta_{j}$, $\eta_{j}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ for $j\neq 0$.
With respect to the first challenge, an important component of DP-CERT is how we handle the privacy concerns of training with augmented data points. We employ augmentation multiplicity, introduced by De et al. (2022) as a method for incorporating standard augmentation methods into DPSGD training. The method involves averaging the gradients of multiple input augmentations of the same training sample before clipping and noising. Since all downstream impact to the model weights from sample $x_{i}$ is contained in this averaged gradient, clipping it provides a finite sensitivity as required for the sampled Gaussian mechanism used in DPSGD (Mironov et al., 2019), and no additional privacy cost is incurred. We propose to apply a slightly modified version of this method for improving CR with DPSGD, namely, by leveraging augmentation multiplicity to train the network with Gaussian noised input data, akin to classical methods of regularizing models for CR (Cohen et al., 2019).
The model updates can be expressed as follows
$\theta_{t+1}=\theta_{t}-\lambda_{t}\left[\frac{1}{|B_{t}|}\sum_{i\in B_{t}}\textrm{clip}_{C}\left(\frac{1}{K+1}\sum_{j=0}^{K}\nabla_{\theta}L(\theta_{t};x_{i}^{j},y_{i})\right)+\frac{\rho C}{|B_{t}|}\xi\right].$ (2)
$\theta_{t}$ denotes the model parameters at iteration $t$, $\lambda_{t}$ is the learning rate, $B_{t}$ is the $t$’th batch, $C$ is the clipping bound, $K$ is the number of augmentations, $\rho$ is the noise multiplier, $\xi\sim\mathcal{N}(0,\textbf{I})$, and $\nabla_{\theta}L(\theta_{t};x_{i}^{j},y_{i})$ is the gradient of the loss for data point $(x_{i}^{j},y_{i})$. Note that $j$ starts from 0, which means we include the original samples along with the augmented ones in model training. We discovered that including the original samples empirically improves the robust accuracy.
Adversarial Training.
When not using Gaussian augmentations to achieve better certified accuracy, we incorporate adversarial training by deploying existing attacks to create adversarial examples. Specifically, we integrate SmoothAdv (Salman et al., 2019) into private training, which, given original data $(x,y)$, optimizes
$\operatorname*{arg\,max}_{\|x^{\prime}-x\|_{2}\leq\epsilon}\bigg{(}-\log\underset{\eta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{E}}[F_{\theta}(x^{\prime}+\eta)_{y}]\bigg{)},$ (3)
to find an $x^{\prime}$ $\epsilon$-close to $x$ that maximizes the cross entropy between the smoothed classifier $g_{\theta}(x^{\prime})$, and label $y$. Using Monte Carlo sampling, Objective (3) can be optimized by iteratively computing the approximate gradient
$\nabla_{x^{\prime}}\bigg{(}-\log\bigg{(}\frac{1}{K}\sum_{j=1}^{K}F_{\theta}(x^{\prime}+\eta_{j})_{y}\bigg{)}\bigg{)}.$ (4)
where $\eta_{1},...,\eta_{K}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. The approximate gradient is then used to update $x^{\prime}$, with the final $x^{\prime}$ used as examples within augmentation multiplicity.
Regularization.
To address the second challenge, the second step in Figure 1 adapts stability regularization to private training in order to minimize the distance between the output probability of the original and augmented examples, thereby improving the robustness to input noise. Stability training (Li et al., 2019) adds a smoothed cross-entropy loss as regularization. Inspired by TRADES (Zhang et al., 2019), we instead use the Kullback–Leibler (KL) divergence with a hyperparameter $\gamma$ controlling the strength of the regularization as:
$L_{\text{stab}}(\theta;x_{i},y_{i})=\sum_{j}L(\theta;x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}F_{\theta}(x_{i})\,\|\,F_{\theta}(x_{i}^{j})\big{)}.$ (5)
Additionally, we propose integrating MACER (Zhai et al., 2020), an alternative training modification to directly optimize the certified accuracy at larger robustness radii without requiring the costly process of adversarial training. MACER achieves this by decomposing the error of a smoothed classifier into a classification error term and a robustness error term, the latter reflecting whether or not the smoothed classifier is able to certify robustness for a given radius. MACER is further described in Appendix B.1.3.
Algorithm 2 DP-CERT Training
0:  Private training set $D=\{(x_{i},y_{i})\mid i\in[N_{\text{prv}}]\}$, loss function $L(\theta_{t};x,y)$, Parameters: learning rate $\lambda_{t}$, noise scale $\rho$, group size $B$, gradient norm bound $C$, number of augmentations $K$.
1:  Initialize ${\theta}_{0}$ randomly
2:  for $t\in[T]$ do
3:     Sample mini-batch $B_{t}$ with sampling probability $B/N_{\text{prv}}$Poisson sampling}
4:     For each $i\in B_{t}$, generate $K$ augmented samples $x_{i}^{j}$Adversarial or Gaussian augmentations}
5:     $L^{\prime}(\theta_{t};x_{i},y_{i})\leftarrow\frac{1}{K+1}\sum_{j=0}^{K}L(\theta_{t};x_{i}^{j},y_{i})+R(\theta_{t},x_{i}^{j})$Add regularizer, average over augmentations}
6:     ${\mathbf{g}}_{t}(x_{i})\leftarrow\nabla_{\theta}L^{\prime}({\theta}_{t};x_{i},y_{i})$Per-sample gradients contain all information from $x_{i}$ and its augmentations}
7:     $\bar{{\mathbf{g}}}_{t}(x_{i})\leftarrow{\mathbf{g}}_{t}(x_{i})/\max\left(1,\frac{\|{\mathbf{g}}_{t}(x_{i})\|_{2}}{C}\right)$Clip gradients}
8:     $\tilde{{\mathbf{g}}}_{t}\leftarrow\frac{1}{|B_{t}|}\left(\sum_{i}\bar{{\mathbf{g}}}_{t}(x_{i})+\mathcal{N}\left(0,\rho^{2}C^{2}{\mathbf{I}}\right)\right)$Add noise to aggregated gradient}
9:      ${\theta}_{t+1}\leftarrow{\theta}_{t}-\lambda_{t}\tilde{{\mathbf{g}}}_{t}$Gradient descent}
10:  end for
11:  Output ${\theta}_{T}$ and compute the overall privacy cost $(\varepsilon,\delta)$ using a privacy accounting method.
Summary of Method.
Table 1: Instantiations of DP-CERT
Method Augmentation Regularization
DP-Gaussian Gaussian None
DP-SmoothAdv Adversarial None
DP-Stability Gaussian Stability
DP-MACER Gaussian MACER
DP-CERT (Figure 1, Algorithm 2) combines augmentation multiplicity, which was originally introduced to improve the accuracy of models trained by DPSGD, with randomized smoothing for inference, and a variety of adversarial augmentation and regularization approaches, which have been shown to improve certified robustness. Comparing the training algorithms 1 and 2 shows that our method follows the template of DPSGD with the following modifications: adversarial or Gaussian perturbations are generated for each data point in the batch; regularization is optionally applied to the loss function; and augmentation multiplicity is used to compute “per-sample” gradients as the average over all augmented data points associated to $x_{i}$. Randomized smoothing at inference time allows certifying robustness, but improved robustness is a result of the training-time modifications.
While DP-CERT is a combination of known methods we emphasize its simplicity compared to previous works that bring CR guarantees to DP. Decreasing complexity is vital for the widespread adoption of robust and private AI, and we simultaneously report improved performance.
Privacy Guarantee.
As described above, augmentation multiplicity achieves the same privacy guarantees as DPSGD by averaging over all augmentations before clipping gradients. This remains true whether random or adversarial augmentations are used, and for any regularized loss function. Once a model is trained with a DP guarantee, it can be used for inference with randomized smoothing at no further privacy cost, thanks to the postprocessing guarantee of DP (Dwork & Roth, 2014; Abadi et al., 2016). Hence, DP-CERT requires no special accounting, and is easily implemented within any mature code package for DP. We note that Gaussian augmentations are a form of input perturbation and may contribute to stronger privacy guarantees. However, input perturbation would first require data points to be clipped for finite sensitivity, impacting performance. Also, we keep the original datapoint that does not have noise added, which is beneficial to performance but not aligned with input perturbation. Hence, we do not attempt to account for DP guarantees from this source of randomness.
In the following sections, we experimentally benchmark four instantiations of the DP-CERT framework as summarized in Table 1.
4 Conclusion
We achieve better certified robustness with DPSGD training through augmentations, regularization, and randomized smoothing, reconciling two crucial objectives for trustworthy ML, namely privacy and robustness. We rely on DP training with augmentations that does not incur additional privacy costs, while employing various regularizations, and adversarial training methods to enhance robustness. Our resulting DP-CERT framework is modular and supports multiple combinations of these methods. Through our extensive experimental study, we confirm that DPSGD training alone, even with state-of-the-art convergence, does not provide satisfactory certified robustness. However, introducing a small number of computationally inexpensive augmentations into training by adding Gaussian noise suffices to yield strong privacy protection and certified robustness, even surpassing much more complex prior methods. By thoroughly analyzing per-sample metrics, we show that the certified radius correlates with the local Lipschitz constant and smoothness of the loss surface; this opens a new path to diagnosing when private models will fail to be robust. In conclusion, our proposed method greatly simplifies the existing solutions to simultaneously achieve CR and DP, and our practical recommendations provide a valuable contribution toward trustworthy ML. When training from scratch, Gaussian augmentations (not adversarial) should be used with DPSGD, and randomized smoothing applied at inference time. For fine-tuning pretrained models, adding stability regularization also helps accuracy, and leads to much lower local Lipschitz constants.
Acknowledgments.
DG, FB, and NP would like to acknowledge sponsors who support their research with financial and in-kind contributions: CIFAR through the Canada CIFAR AI Chair, NSERC through a Discovery Grant, the Ontario Early Researcher Award, and the Sloan Foundation. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Broader Impact Statement
Nowadays, machine learning finds extensive application in our society. As a result, ensuring the integrity of the models we build is vital. This entails safeguarding individuals’ privacy while maintaining the models’ robustness. Our paper focuses on examining the intersection of privacy and robustness for machine learning models, both of which are essential for establishing trustworthy ML. We introduce the DP-CERT framework, which, compared to prior work, simplifies the process for machine learning developers to create models that are both robust and private, while increasing the robustness afforded for a given privacy level.
While both differential privacy and certified robustness can provide mathematical guarantees on the privacy and robustness of trained models, the guarantees are probabilistic and can randomly fail. For DP, failures roughly occur with probability $\delta$, while for CR, some specific datapoints may not be certifiable for adequately large radii $r$. Practitioners should evaluate their risk tolerances for failures of privacy and robustness, and choose the parameters of these guarantees appropriately.
Furthermore, it has been noted that differential privacy can have particularly negative impacts on the fairness of models (Bagdasaryan et al., 2019; Xu et al., 2021; Esipova et al., 2023). While we do not specifically address fairness aspects in this paper, it is another prong of trustworthy ML that should be taken into account and ensured for any ML model that is to be deployed.
Appendix A Differential Privacy Background
A.1 Differential Privacy and DPSGD
Differential Privacy (DP) (Dwork et al., 2006) is a formal framework that aims to provide mathematical guarantees on the privacy of individual data points in a dataset, while allowing one to learn properties over the entire dataset. More formally a randomized mechanism $M$ fulfills $(\varepsilon,\delta)$-DP if, for all pairs $D$ and $D^{\prime}$ of neighbouring datasets (i.e. datasets differing in only one data point) and all sets $S$ of output values of $M$,
$\mathrm{Pr}\left[M(D)\in S\right]\leq e^{\varepsilon}\mathrm{Pr}\left[M(D^{\prime})\in S\right]+\delta,$ (6)
where $\varepsilon$ and $\delta$ are privacy parameters. $\varepsilon$ indicates the level of obtained privacy, with a smaller value of $\varepsilon$ indicating a stronger privacy guarantee. $\delta$ accounts for the probability that the algorithm will suffer a catastrophic privacy failure, which occurs when the privacy loss random variable becomes infinite (Kifer et al., 2022). This will happen when an event with zero probability for one dataset has non-zero probability for a neighbouring dataset, meaning that there is no possible multiplicative factor $e^{\varepsilon}$ that could satisfy the DP inequality Eq. 6 with $\delta=0$. Hence, a larger value of $\delta$ increases the possibility of unmitigated privacy leakage.
The most popular algorithm for implementing DP guarantees in ML is differentially private stochastic gradient descent (DPSGD) (Abadi et al., 2016). It extends the standard SGD algorithm with two additional steps, namely gradient clipping and noise addition. While the former bounds the sensitivity of the model update, the latter implements the privacy guarantee by preventing the gradients from revealing too much information about individual data points. More formally, let $\theta_{t}$ denote the model parameters at training iteration $t$. At each iteration $t$, DPSGD computes the gradient of the loss function with respect to $\theta_{t}$ at an individual data point $x$ as $\nabla_{\theta}L(\theta_{t},x)$. The data point’s gradient is then clipped to a maximum norm $C$ using the operation $\text{clip}\left(\nabla_{\theta}L(\theta_{t},x),C\right)$, which replaces the gradient with a vector of the same direction but smaller magnitude if its norm exceeds $C$. The clipped gradient from all datapoints in a batch are aggregated, then perturbed by adding random noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$, where $\rho$ is the noise scale parameter. We detail DPSGD in Algorithm 1.
A.2 PSAC
Per-sample adaptive clipping (PSAC) (Xia et al., 2023) is one of a number of approaches that tries to reduce the bias from per-example clipping (Bu et al., 2023a; b; Esipova et al., 2023), and was motivated for maximizing the signal to noise ratio of gradient updates. It has shown the best performance for differentially private models on several datasets including MNIST, Fashion-MNIST and CIFAR10. To complement DPSGD as an algorithm with better convergence, we incorporated PSAC into our experiments.
PSAC is similar to DPSGD, with the exception that it employs a different clipping method,
$\textrm{clip}_{C,r}(g_{i,t})=C\cdot g_{i,t}\bigg{/}\biggl{(}\|g_{i,t}\|+\frac{r}{\|g_{i,t}\|+r}\biggr{)}.$ (7)
Where $g_{i,t}$ denotes the loss gradient for the $i$th sample at iteration $t$. The motivation of this clipping method is that per-example gradients with small norms come from data points on which the model has already converged, and these gradients are often orthogonal to mini-batch gradients. Hence, small norm gradients should not have disproportionally large contributions to the batch gradient as when using per-sample normalization methods (Bu et al., 2023b). Compared to such approaches, PSAC reduces the influence of small norm gradients, and empirically shows better convergence and utility.
Appendix B Additional Background and Related Work
B.1 Randomized Smoothing
Previous works have tackled improving the certified robustness of randomized smoothing methods in a variety of ways. The dominant approach for doing so involves modifications to training the base classifier so as to increase robustness and accuracy under Gaussian perturbations. The simplest approach involves adding noise to inputs during training (Cohen et al., 2019), while other works utilize regularization (Zhai et al., 2020; Li et al., 2019; Jeong & Shin, 2020), ensembling (Horváth et al., 2022), and adversarial training (Salman et al., 2019). While these modifications have been independently studied in the context of improving the certified accuracy of randomized smoothing classifiers, our work is the first to integrate these methods with private training through augmentation multiplicity. We provide additional information on two of these training modification methods mentioned in Section 3, as well as Consistency regularization used in one ablation experiment.
B.1.1 SmoothAdv
One of the most effective methods for improving the performance of a randomized smoothing classifier utilizes adversarial training of the classifier. The method, SmoothAdv proposed in (Salman et al., 2019), was motivated by the idea that to improve certified accuracy at a larger certification radius one needs a classifier that is more robust to local perturbations, and the best known method of achieving that is through adversarial training.
Given a soft classifier $F:\mathbb{R}^{d}\rightarrow P(Y)$ where $P(Y)$ is the set of probability distributions over $Y$, its smoothed soft classifier $G$ is defined as:
$G(x)=(F*\mathcal{N}(0,\sigma^{2}I))(x)=\mathbb{E}_{\eta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x+\eta)].$ (8)
The goal of SmoothAdv is to find a point $\hat{x}$ that maximizes the cross entropy loss of $G$ in an $l_{2}$ ball around $x$. SmoothAdv uses a projected gradient descent variant to approximately find $\hat{x}$, and define $J(x^{\prime})=L_{\textrm{CE}}(G(x^{\prime}),y)$ to compute
$\nabla_{x^{\prime}}J(x^{\prime})=\nabla_{x^{\prime}}\biggl{(}-\log\mathbb{E}_{\eta\sim\mathcal{N}(0,\sigma^{2}I)}[F(x^{\prime}+\eta)_{y}]\biggr{)}.$ (9)
Since the expectation in Equation 9 is difficult to compute exactly, a Monte Carlo approximation is used by sampling noise $\eta_{1},...,\eta_{m}\sim\mathcal{N}(0,\sigma^{2}I)$ to approximately compute $\nabla_{x^{\prime}}J(x^{\prime})$,
$\nabla_{x^{\prime}}J(x^{\prime})\approx\nabla_{x^{\prime}}\biggl{(}-\log\biggl{(}\frac{1}{m}\sum\limits_{i=1}\limits^{m}[F(x^{\prime}+\eta_{i})_{y}]\biggr{)}\biggr{)}.$ (10)
Finally, $x^{\prime}$ is updated by taking a step in the direction of $\nabla_{x^{\prime}}J(x^{\prime})$, and the final $x^{\prime}$ is used to train the classifier.
B.1.2 Consistency Regularization
For ablation purposes, we also compare to Consistency regularization (Jeong & Shin, 2020), a method very similar to Stability training which instead minimizes the KL divergence between $\hat{F}_{\theta}(x_{i})$ and $F_{\theta}(x_{i})$, where $\hat{F}_{\theta}(x_{i})=\frac{1}{K}\sum_{j}F_{\theta}(x_{i}^{j})$ is the average output probability of all smoothed samples. The loss can be expressed as
$\ L_{\text{cons}}(x_{i},y_{i})=\sum_{j}L_{\textrm{CE}}(x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}\hat{F}_{\theta}(x_{i})||F_{\theta}(x_{i}^{j})\big{)}.$ (11)
B.1.3 MACER
Similar to the Stability training method, MACER (Zhai et al., 2020) also modifies the loss for optimization so that the final model has higher certified accuracy at a larger certified radius. In contrast to SmoothAdv, regularizing the model to be more robust in this sense does not require generation of adversarial examples and instead can be optimized directly. The method achieves this by decomposing the error of the smoothed classifier into a classification error term and a robustness error term. The former captures the error from the smoothed classifier misclassifying a given datapoint and the latter captures the error of a certified radius being too small.
Robustness error, much like hard label classification error, cannot be optimized directly. To address this, MACER proposes a surrogate loss to minimize the robustness error term—a hinge loss on the data $(x,y)$ for which $g_{\theta}(x)=y$,
$\max\{0,\gamma-(\Phi^{-1}(\hat{f}_{\theta}(x)_{y})-\Phi^{-1}(\hat{f}_{\theta}(x)_{\hat{y}\neq y}))\},$ (12)
where $\hat{f}_{\theta}(x)$ denotes the average of softmax probabilties on Gaussian perturbations of $x$, $\hat{f}_{\theta}(x)_{y}$ denotes the softmax probability $\hat{f}_{\theta}(x)$ assigned to the true class $y$, and $\hat{f}_{\theta}(x)_{\hat{y}\neq y}$ denotes the maximum softmax probability over classes other than the true class. This loss term is added as a regularization term to the cross entropy loss from the soft smoothed classifier.
B.2 DPSGD and Robustness
The adversarial robustness of differentially private models has been studied in several prior works. Tursynbek et al. (2021) demonstrated that models trained with DPSGD are sometimes more vulnerable to input perturbations. Boenisch et al. (2021) further consolidate this claim with more experiments, and showed that improper choices of hyperparameters can lead to gradient masking. Zhang & Bu (2022) found that the success of adversarially training robust models with DPSGD depends greatly on choices of hyperparameters, namely smaller clipping thresholds and learning rates, differing from those that produce the most accurate models. These works reveal interesting adversarial robustness characteristics of DP models, however, they do not endeavor to improve the robustness of DP models.
Bu et al. (2022a) propose combining adversarial training with DPSGD. They achieve this by replacing the original example with an adversarially crafted example, obtained with a FGSM or PGD attack. Their work is orthogonal to ours in that they focus on the robustness to adversarial attacks, whereas we focus on certified robustness. Additionally, they do not perform any data augmentation, which has been shown to be effective against adversarial attacks (Rebuffi et al., 2021).
Compared with existing works in DP and CR (Phan et al., 2019; 2020; Tang et al., 2022) reviewed in Section 2, DP-CERT 1) presents a simple and effective training scheme with augmentations, allowing practitioners to introduce different adversarial training techniques through noising, regularization, and adversarially crafted examples, 2) does not rely on a denoiser at inference time, reducing inference latency, and 3) can be used for training both randomly initialized or pre-trained networks.
Appendix C Experimental Details
C.1 Datasets
We conduct experiments on MNIST, Fashion-MNIST and CIFAR10. The MNIST database of handwritten digits has a training set of 60,000 examples, and a test set of 10,000 examples, as does Fashion-MNIST. Each example in MNIST and Fashion-MNIST is a 28$\times$28 grayscale image, associated with one label from 10 classes.
The CIFAR10 dataset consists of 60,000 RGB images from 10 classes, with 6,000 images per class. There are 50,000 training images and 10,000 test images of size 32$\times$32$\times$3.
C.2 Code and Implementation
Our implementation of DP-CERT in code is provided as supplementary material. We give credit to the original repository for the implementation of SmoothAdv, MACER, Stability, and Consistency regularization. [^1] https://github.com/jh-jeong/smoothing-consistency For CERTIFY [^2] https://github.com/locuslab/smoothing and computing local Lipschitz constants [^3] https://github.com/yangarbiter/robust-local-lipschitz, we use the code provided by the original authors. All experiments were conducted on a cluster of 8 Nvidia V100 GPUs. All the training and inference procedures are implemented based on Pytorch v1.13.0 (Paszke et al., 2019) and Opacus v1.3.0 (Yousefpour et al., 2021), and our code is provided at github.com/layer6ai-labs/dp-cert.
C.3 Additional Implementation and Hyperparameters
We set the learning rate as 0.001 and train the models for 10 epochs. The rest of the hyperparameters are the same as used by Bu et al. (2022b). For evaluation, we use CERTIFY with parameters $n=10,000$, $n_{0}=100$, and $\alpha=0.001$, following previous work (Cohen et al., 2019; Salman et al., 2019). Our implementations of DP-SmoothAdv, DP-Stability and DP-MACER are adapted from the original codebases for those regularization approaches (Salman et al., 2020; Li et al., 2019; Zhai et al., 2020), and we use the default hyperparameters reported in the original papers. We set the number of augmentations $K$ to 2 for MNIST and Fashion-MNIST, and 1 for CIFAR10, as they bring a better trade-off between certified accuracy and efficiency (see Figure LABEL:fig:ablation).
Figure 2: Approximate certified accuracy comparison on MNIST (top), and Fashion-MNIST (bottom).
Appendix D Additional Results
In this section, we show the complete results and figures for our comparative study, ablations, and analysis of fine-grained metrics.
D.1 Additional Comparative Study
Figure 2 complements Figure LABEL:fig:certified_accuracy_main from the main text by comparing the approximate certified accuracy between DP-CERT and baselines on MNIST, and Fashion-MNIST. The results confirm that Regular and DP training methods that do not specifically consider robustness as a criteria give poor CR results, but all instantiations of the DP-CERT framework do much better. We are unable to show further results similar to Figure LABEL:fig:certified_accuracy_transdenoiser using the baselines SecureSGD, StoBatch, and TransDenoiser because their implementations are highly customized and not compatible with our codebase based on Pytorch and Opacus. The baseline results in Figure LABEL:fig:certified_accuracy_transdenoiser are directly copied from (Tang et al., 2022) which provided limited experimental settings for us to match.
Figure 3: Certified accuracy of DP-Gaussian and DP-SmoothAdv combined with consistency regularization and PSAC on MNIST (top) and Fashion-MNIST (bottom).
Figure 4: Certified accuracy comparison under different numbers of augmentations (0, 2, 4, 6 and 8) on MNIST (top) and Fashion-MNIST (bottom). The line marked 0 indicates replacing the original datapoint with an augmentation.
D.2 Additional Ablation Study
To extend Figure LABEL:fig:ablation from the main text, we complete the ablation study in Figures 3 and 4 on MNIST and FashionMNIST under $\sigma\in\{0.25,0.5,1.0\}$. Figure 3 shows the effects of adding Consistency regularization, or changing the DP clipping method to PSAC. Neither modification leads to significant changes to the certified accuracy, so we recommend using the simplest combination of standard DPSGD clipping without additional regularization when training models from scratch. Figure 4 shows the effects of changing the multiplicity of augmentations $K$. Since additional augmentations greatly increase computation time, we must navigate the tradeoff between efficiency and performance. We see that it is beneficial to use some number of augmentations or replace the original datapoint with a single augmentation (marked line 0), but beyond $K=2$ there is little further gain in certified accuracy. We note that the case without augmentations at all, namely DPSGD, achieves much worse CR as presented in Table LABEL:tab:certified_accuracy_main and Figures LABEL:fig:certified_accuracy_main and 2. One of our main conclusions is that adding a small number of Gaussian augmentations to DPSGD is sufficient to improve certified robustness at a low cost to efficiency.
Table 2 and 3 show additional hyperparameter studies on learning rate $\lambda$ and clipping bound $C$ for DP-Gaussian under $\sigma=0.25$. Not only does this show that our models were well-tuned for clean accuracy, but that there is no dramatic drop in performance for nearby parameter values.
Table 2: Effect of Learning Rate $\lambda$ on Clean Accuracy (%)
Dataset Learning Rate $\lambda$ Clean Accuracy (%)
MNIST 0.25 97.71
0.5 98.13
1 97.78
Fashion-MNIST 2 83.47
4 84.76
8 83.96
CIFAR10 0.0005 84.11
0.001 87.61
0.002 85.53
Table 3: Effect of Clipping Norm $C$ on Clean Accuracy (%)
Clipping Norm $C$ Clean Accuracy (%)
MNIST Fashion-MNIST CIFAR10
0.01 94.89 75.60 83.79
0.05 97.75 83.42 83.99
0.1 98.13 84.76 87.61
0.2 97.72 84.00 84.19
0.5 96.47 81.46 84.24
D.3 Additional Per-sample Metric Analysis
Figures 5 through 9 complete our presentation of the three proposed metrics for interpreting robustness, the input gradient norms, input Hessian spectral norms, and local Lipschitz constants, over three datasets and values of $\sigma$ (see also Figure LABEL:fig:analysis-mnist-0.5 in the main text). We use various training methods on MNIST and Fashion-MNIST under $\sigma\in\{0.25,0.5,1.0\}$. Echoing the analysis of RQ1 in Section LABEL:sec:fine_grained_analysis, DPSGD produces a bimodal distribution for the Hessian and gradient norms, while Regular training exhibits a log-normal distribution and smaller tails for large metric values. PSAC shifts the distributions to be closer to those of Regular training by reducing clipping bias. DP-CERT methods, on the other hand, shift the distribution towards smaller metric values, resulting in higher certified accuracy. An exception is DP-Stability, which has significantly higher average gradient and Hessian norms, but without the mode at very high values, and with lower local Lipschitz constants than the other three variants.
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 5: Per-sample metric comparison, MNIST, $\sigma=0.25$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 6: Per-sample metric comparison, MNIST, $\sigma=1.0$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 7: Per-sample metric comparison, Fashion-MNIST, $\sigma=0.25$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 8: Per-sample metric comparison, Fashion-MNIST, $\sigma=0.5$
(a) Comparison of Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Comparison of DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 9: Per-sample metric comparison, Fashion-MNIST, $\sigma=1.0$
Figure 10 shows the distribution of certified radii for baseline training methods and an instance of DP-CERT for the same settings as Figure LABEL:fig:analysis-mnist-0.5. Whereas Regular, DPSGD, and PSAC training all have a large spike of samples that cannot be certified at any level, DP-Gaussian achieves certified radii above 1.0 for most samples, with the mode even higher at 1.6.
Figure 10: Distributions of certified radii for Regular, DPSGD, PSAC, and DP-Gaussian training on MNIST with $\sigma=0.5$.
Figure 11: Adversarial accuracy against a $L_{\infty}$-FGSM attack under various attack strengths on MNIST (top) and Fashion-MNIST (bottom).
While our main focus has been on certified robustness, we briefly examine adversarial robustness. Figure 11 shows the adversarial accuracy under a $l_{\infty}$-FGSM attack, with the attack strength in $\{0.0005,0.01,0.1,0.5,1\}$. Consistent with the ranking of the average local Lipschitz constant from Figures 5 through 9, DP-Stability consistently outperforms other approaches, while DP-Gaussian, DP-SmoothAdv, and DP-MACER all achieve similar adversarial accuracy above that of the unprotected baselines.
(a) Per-sample metric comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Per-sample metric comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 12: Distributions of log local Lipschitz constants among baselines and proposed methods, MNIST, $\sigma=0.5$. In each subfigure, data points are classified into ones with certified radius above the threshold $\tau$ (top row), and below the threshold (bottom row). We display $\tau\in\{0.5,1.0,1.5\}$.
(a) Per-sample metric comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Per-sample metric comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 13: Comparing the distribution of input gradient norm of baselines and proposed methods on MNIST. All DP-CERT methods are trained with $\sigma=0.5$. We show the logarithmic metric values for all methods for better visualization.
Finally, as shown in Figure 12, we select three certified radius thresholds $\tau\in\{0.5,1.0,1.5\}$, and separately plot the log of local Lipschitz constants of examples above and below $\tau$ on the top and bottom rows of the subfigures respectively. First, the examples with certified radii below the threshold have higher average local Lipschitz constant. Second, as we increase the threshold $\tau$, more examples with higher local Lipschitz constant end up below the certified radius threshold. We do a similar analysis for the other two metrics in Figures 13 and 14, but since the local Lipschitz constant is derived using a PGD attack, it naturally correlates with the robustness to adversarial examples better. For further comparison, we present the FGSM accuracy on MNIST and Fashion-MNIST under attack strengths in $\{0.0005,0.01,0.1,0.5,1\}$ in Figure 11. Consistent with the ranking of the average local Lipschitz constant, DP-Stability consistently outperforms other approaches, while DP-Gaussian, DP-SmoothAdv and DP-MACER achieve similar adversarial accuracy over different attack margins.
(a) Per-sample metric comparison among Regular, DPSGD, PSAC, and DP-Gaussian training.
(b) Per-sample metric comparison among DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER.
Figure 14: Comparing the distribution of input Hessian spectral norm of baselines and proposed methods on MNIST. All DP-CERT methods are trained with $\sigma=0.5$. We show the logarithmic metric values for all methods for better visualization.